OpenStudy (anonymous):

for f(x)= [sin(x)]/x find f'(x) and f'(c) when c= pi/6 I am stuck on the second part!

OpenStudy (anonymous):

Ok, so I found the first part: f'(x)= [(cosx)x - (sinx)]/x^2 So for the second part, I have to put pi/6 in: f'(pi/6)= [(cospi/6)pi/6 - sinpi/6]/(pi/6)^2 But I am not sure what to do. I could solve with my calculator but I think my teacher wants it in radian form. Can someone point me in the right direction?

myininaya (myininaya):

Do you know the unit circle?

OpenStudy (anonymous):

Not by heart, but I have a picture on my computer. I'll pull it up.

myininaya (myininaya):

Use that to evaluate cos(pi/6) and sin(pi/6)

OpenStudy (anonymous):

Would it be sqrt3/2?

myininaya (myininaya):

That is for cos(pi/6) and sin(pi/6)=1/2 right?

myininaya (myininaya):

\[f'(\frac{\pi}{6})=\frac{\frac{\sqrt{3}}{2} \cdot \frac{\pi}{6}-\frac{1}{2}}{(\frac{\pi}{6})^2}\] So this is what you have. Most teachers do not like compound fractions. Do you know how to clear the compound fractions?

OpenStudy (anonymous):

Right, so sqrt3/2(pi/6) -1/2 all over pi/6^2 So what next?

myininaya (myininaya):

look at the bottoms of the mini-fractions. The highest number in them is the 36, but all the denominators of the mini-fractions go into 36 so multiply top and bottom of the big fraction by 36.

OpenStudy (anonymous):

36? Sorry, but I'm a bit lost. Where did you get that from?

myininaya (myininaya):

The 36 on bottom of the mini-fraction in the big fraction.

myininaya (myininaya):